# A Closer Look at Time Steps is Worthy of Triple Speed-Up for Diffusion Model Training

### v1

@@ 3.4 Compatibility with other acceleration methods
Until now, we evaluate the effectiveness and generalization of our proposed method: SpeeD is a task-agnostic and architecture-agnostic diffusion acceleration approach. Is SpeeD compatible with other acceleration techniques? To investigate this, we apply our approach with two recent proposed algorithms: masked diffusion transformer (MDT) gao2023masked and fast diffusion model (FDM) (wu2023fast, ).
MDT + SpeeD. MDT (gao2023masked, ) proposes a masked diffusion transformer method, which applies a masking scheme in latent space to enhance the contextual learning ability of diffusion probabilistic models explicitly. MDT can speed up the diffusion training by 10 times. They evaluate their MDT with DiT-S/2. We just inject our SpeeD on their MDT and report FID-Iteration curves for comparison in Fig. 5(a). All the results are obtained on ImageNet-1K dataset. Our approach can further accelerate MDT at least by 4 $\times$, which indicates the good compatibility of SpeeD.
FDM + SpeeD. Fast Diffusion Model wu2023fast is a diffusion process acceleration method inspired by the classic momentum approach to solve optimization problem in parameter space. By integrating momentum into the diffusion process, it achieves similar performance as EDM (karras2022elucidating, ) with less training cost. Based on the official implementation, we compare EDM, FDM, and FDM + SpeeD on CIFAR-10 of $32\times 32$ images. FDM accelerates the EDM by about 1.6 $\times$. Meanwhile, SpeeD can further reduce the overall training cost around 1.6 $\times$.
@@ 3.5 Ablation Experiments
(a) Components of SpeeD. Suppressing some trivial time steps do help.
sampling curve CAW FID $\downarrow$ uniform 17.37 uniform 16.75 asymmetric 15.82 asymmetric 15.07
@@
suppression intensity $k$ FID $\downarrow$ 1 15.01 5 14.86 10 16.97 25 25.59
(c) Symmetry ceiling $\lambda$. Weighting should be in moderation.
symmetry ceiling $\lambda$ FID $\downarrow$ 0.5 15.46 0.6 14.86 0.8 16.83 1.0 23.77
@@
3.6 Visualization
(a) Visualization comparison on MetFaces dataset.
(b) Visualization comparison on FFHQ dataset.
Figure 6: SpeeD obtains significant improvements than baseline in visualizations. More visualizations on other datasets and tasks can be found in the Appendix D.
@@ 4 Related Work
Diffusion models have emerged as the dominant approach in generative tasks (saharia2022photorealistic, ; chen2023humanmac, ; wang2024neural, ; igashov2024equivariant, ), which outperform other generative methods including GANs (zhu2017unpaired, ; isola2017image, ; brock2018large, ), VAEs (kingma2013auto, ), flow-based models (dinh2014nice, ). These methods (ho2020denoising, ; song2020denoising, ; karras2022elucidating, ) are based on non-equilibrium thermodynamics (jarzynski1997equilibrium, ; sohl2015deep, ), where the generative process is modeled as a reverse diffusion process that gradually constructs the sample from a noise distribution (sohl2015deep, ). Previous works focused on enhancing diffusion models’ generation quality and alignment with users in visual generation. To generate high-resolution images, Latent Diffusion Models (LDMs) (saharia2022photorealistic, ; rombach2022high, ) perform diffusion process in latent space instead of pixel space, which employ VAEs to be encoder and decoder for latent representations.
@@
To reduce the computational costs, previous works accelerate the diffusion models in training and inference. For training acceleration, the earliest works (choi2022perception, ; hang2023efficient, ) assign different weights to each time step on Mean Square Error (MSE) loss to improve the learning efficiency. The other family of methods in training acceleration are from different perspectives, e.g., network architecture (ryu2022pyramidal, ; wang2024patch, ) and diffusion algorithm (karras2022elucidating, ; wu2023fast, ). Most recently, masking modeling diffusion models (gao2023masked, ; zheng2023fast, ) are been proposed to train diffusion models with reconstructing tasks on masked tokens. SpeeD is a general method to enhance the training process tailored for diffusion models with good compatibility with these methods. In another field of acceleration, there are a number of works focus on improving the sampling speed, due to the slow inference speed of diffusion models. The inference is slow because diffusion models transit the noise into target distribution gradually. Methods are proposed to speed up by using fewer reverse steps while maintaining sample quality, including DDIM (song2020denoising, ), Analytic-DPM (bao2022analytic, ), and DPM-Solver (lu2022dpm, ).
@@
To better control the generation process with various conditions, e.g., image style, text prompt and stroke, Classifier-Free Guidance (CFG) proposes a guidance strategy with diffusion models that balance the sample quality and prompt alignment. ControlNet (zhang2023adding, ) reuses the large-scale pre-trained layers of source models to build a deep and strong encoder to learn specific conditions. Recently benefiting from diffusion models in the image generation field, video generation (ma2024latte, ; pku_yuan_lab_and_tuzhan_ai_etc_2024_10948109, ) is getting trendy.
@@ 5 Conclusion
Limitations and future work. Diffusion training acceleration is largely different from its inference acceleration. The research for training acceleration costs much more computational resources than inference acceleration. Therefore, maybe our experiments are not sufficient. In the future, we will explore the effectiveness of applying our approach to more resource-density diffusion task, such as video generation Sora ; zhang2023show ; jiang2023videobooth . Another direction is to explore the potential in more diffusion algorithms, such as the consistency model song2023consistency . The adoption of SpeeD not only advances the state-of-the-art in diffusion model training efficiency but also aligns with global efforts towards sustainable and eco-friendly AI development. This method demonstrates that high-performance machine learning does not have to come at the cost of environmental degradation.
@@ Appendix A More Detail of Experiments
A.1 Architecture and Training Recipe.
We utilize U-Net and DiT as our base architecture in the diffusion model. pre-trained VAE which loads checkpoints from huggingface is employed to be latent encoder. Following U-Net implementation from LDM and DIT from official implementation, we provide the architecture detail in Tab. 6. We provide our basic training recipe and evaluation setting with specific details in Tab. 7.
@@
Table 6: Architecture detail of U-Net and DiT on MetFaces, FFHQ, and ImageNet.
@@ A.2 Datasets
A.3 Detail of MDT + SpeeD Experiment
@@
Following the official implementation of MDT, We utilize DiT-S/2 and MDT-S/2 as our base architecture, whose total block number both is 12 and the number of decoder layers in MDT is $2$. We employ the AdamW [36] optimizer with constant learning rate 1e-4 using $256$ batch size without weight decay on class-conditional ImageNet with an image resolution of $256^{2}$. We perform training on the class-conditional ImageNet dataset with images of resolution 256x256. The diffusion models are trained for a total of 1000K iterations, utilizing a mask ratio of 0.3.
A.4 Detail of FDM + SpeeD Experiment
FDM add the momentum to the forward diffusion process with a scale that control the weight of momentum for faster convergence to the target distribution. Following official implementation, we train diffusion models of EDM and FDM. We retrain these official network architecture which is U-Net with positional time embedding with dropout rate 0.13 in training. We adopt Adam optimizer with learning rate 1e-3 and batch size 512 to train each model by a total of 200 million images of $32^{2}$ CIFAR-10 dataset. During training, we adopt a learning rate ramp-up duration of 10 Mimgs and set the EMA half-life as 0.5 Mimgs. For evaluation, EMA models generate 50K images using EDM sampler based on Heun’s $2^{nd}$ order method [55].
A.5 Text-to-Image Experiment Detail
@@
A.6 Theoretical Analysis
A.6.1 Notations
@@
A.6.2 Auxiliary Lemma and Core Theorem
@@
In DDPM [19], using a simple equivariant series $\{\beta_{t}\}_{t\in[T]}$ to simplify the complex cumulative products $\{\bar{\alpha}_{t}\}_{t\in[T]}$, we obtain the following auxiliary upper bound of $\bar{alpha}_{t}$.
@@ A.6.4 Proof of Lemma 1
A.6.5 Proof of Theorem 1
Before the proof of the theorem, we note that the samples $x_{t}|x_{0}\sim\mathcal{N}(\mu_{t},\Sigma_{t})$ have the following bounds with Lemma 1:
@@ Appendix B More Experiment Results
Super resolution with SpeeD. We employ SpeeD to super-resolution image generation on 512 $\times$ 512 MetFaces compared with vanilla DiT. We train DiT-XL/2 for 100K training iterations and compare the FID score at 50K, 100K training iterations. The batch size is $32$ for saving the GPU memory. As shown in 8, SpeeD obtain better performance than vanilla DiT at same training iterations on $512^{2}$ MetFaces dataset. It indicates that SpeeD can achieve training acceleration on super-resolution tasks.
Appendix C Other Related Works
@@
We discuss other works related to SpeeD, including Text to Image and Video generation. Another point to mention is that we learn from InfoBatch [43] in writing.
@@
Text to image generation with diffusion models Text-to-image generation has emerged as a hotly contested and rapidly evolving field in recent years, with an explosion of related industrial products springing up [49, 46, 2, 6, 11]. Convert textual descriptions into corresponding visual content, models not only learn to synthesize image content but also ensuring alignment with the accompanying textual descriptions. To better align images with textual prompt guidance, previous work has primarily focused on enhancements in several schemes including strengthening the capacity of text encoder [45, 44] improving the condition plugin module in diffusion model [65], improving data quality [2].
@@
Video generation with diffusion models. As diffusion models achieve tremendous success in image generation, video generation has also experienced significant breakthroughs, marking the field’s evolution and growth. Inspired by image diffusion, pioneering works such as RVD [62] and VDM [20] explore video generation using diffusion methods. Utilizing temporal attention and latent modeling mechanisms, video diffusion has advanced in terms of generation quality, controllability, and efficiency [18, 51, 67, 16, 64, 14, 59, 57]. Notably, Stable Video Diffusion [3] and Sora [5] achieve some of the most appealing results in the field.
@@
Other diffusion acceleration works To achieve better results with fewer NFE steps, Consistency Models [54] and Consistency Trajectory Models [28] employ consistency loss and novel training methods. Rectified Flow [34], followed by Instaflow [35], introduces a new perspective to obtain straight ODE paths with enhanced noise schedule and improved prediction targets, together with the reflow operation.

### v2

@@ 3.5 Compatibility with other acceleration methods
Until now, we evaluate the effectiveness and generalization of our proposed method: SpeeD is a task-agnostic and architecture-agnostic diffusion acceleration approach. Is SpeeD compatible with other acceleration techniques? To investigate this, we apply our approach with two recent proposed algorithms: masked diffusion transformer (MDT) (Gao et al., 2023) and fast diffusion model (FDM) (Wu et al., 2023).
MDT + SpeeD. MDT (Gao et al., 2023) proposes a masked diffusion transformer method, which applies a masking scheme in latent space to enhance the contextual learning ability of diffusion probabilistic models explicitly. MDT can speed up the diffusion training by 10 times. They evaluate their MDT with DiT-S/2. We just inject our SpeeD on their MDT and report FID-Iteration curves for comparison in Fig. 6(a). All the results are obtained on ImageNet-1K dataset. Our approach can further accelerate MDT at least by 4 $\times$, which indicates the good compatibility of SpeeD.
FDM + SpeeD. Fast Diffusion Model (Wu et al., 2023) is a diffusion process acceleration method inspired by the classic momentum approach to solve optimization problem in parameter space. By integrating momentum into the diffusion process, it achieves similar performance as EDM (Karras et al., 2022) with less training cost. Based on the official implementation, we compare EDM, FDM, and FDM + SpeeD on CIFAR-10 of $32\times 32$ images. FDM accelerates the EDM by about 1.6 $\times$. Meanwhile, SpeeD can further reduce the overall training cost around 1.6 $\times$.
@@ 3.6 Ablation Experiments
(a) Components of SpeeD. Suppressing some trivial time steps does help.
sampling curve CAW FID $\downarrow$
uniform 17.37
uniform 16.75
asymmetric 15.82
asymmetric 15.07
@@
suppression intensity $k$ FID $\downarrow$
1 15.01
5 14.86
10 16.97
25 25.59
(c) Symmetry ceiling $\lambda$. Weighting served as temperature factor should be in moderation.
symmetry ceiling $\lambda$ FID $\downarrow$
0.5 15.46
0.6 14.86
0.8 16.83
1.0 23.77
@@ 4 Related Work
Diffusion models have emerged as the dominant approach in generative tasks (Saharia et al., 2022; Chen et al., 2023b; Wang et al., 2024a; Igashov et al., 2024), which outperform other generative methods including GANs (Zhu et al., 2017; Isola et al., 2017; Brock et al., 2018), VAEs (Kingma & Welling, 2013), flow-based models (Dinh et al., 2014). These methods (Ho et al., 2020; Song et al., 2020; Karras et al., 2022) are based on non-equilibrium thermodynamics (Jarzynski, 1997; Sohl-Dickstein et al., 2015), where the generative process is modeled as a reverse diffusion process that gradually constructs the sample from a noise distribution (Sohl-Dickstein et al., 2015). Previous works focused on enhancing diffusion models’ generation quality and alignment with users in visual generation. To generate high-resolution images, Latent Diffusion Models (LDMs) (Saharia et al., 2022; Rombach et al., 2022) perform diffusion process in latent space instead of pixel space, which employ VAEs to be encoder and decoder for latent representations.
@@
To reduce the computational costs, previous works accelerate the diffusion models in training and inference. For training acceleration, the earliest works (Choi et al., 2022; Hang et al., 2023) assign different weights to each time step on Mean Square Error (MSE) loss to improve the learning efficiency. The other methods in training acceleration are proposed, e.g., network architecture (Ryu & Ye, 2022; Wang et al., 2024b) and diffusion algorithm (Karras et al., 2022; Wu et al., 2023). Masking modeling (Gao et al., 2023; Zheng et al., 2023) are recently proposed to train diffusion models. Works  (Go et al., 2023; Park et al., 2024b; a; Kim et al., 2024) provide observations for explanation from the perspective of multi-tasks learning. SpeeD is of good compatibility with these methods, e.g., (Yu et al., 2024; Zheng et al., 2024a; b). In the field of sampling acceleration, a number of works tackle the slow inference speed of diffusion models by using fewer reverse steps while maintaining sample quality, including DDIM (Song et al., 2020), Analytic-DPM (Bao et al., 2022), and DPM-Solver (Lu et al., 2022).
@@
To better control the generation process with various conditions, e.g., image style, text prompt and stroke, Classifier-Free Guidance (CFG) proposes a guidance strategy with diffusion models that balance the sample quality and prompt alignment. ControlNet (Zhang et al., 2023b) reuses the large-scale pre-trained layers of source models to build a deep and strong encoder to learn specific conditions. Recently benefiting from diffusion models in the image generation field, video generation (Ma et al., 2024; Lab & etc., 2024) is getting trendy.
@@ Appendix A More Detail of Experiments
A.1 Statement
Experimental design note: Referring to previous works (Peebles & Xie, 2023; Go et al., 2023; Park et al., 2024b; Choi et al., 2022; Hang et al., 2023; Esser et al., 2024; Xu et al., 2024), our experiments train 50K, 100K, and 400K in MetFaces 256 × 256, FFHQ 256 × 256, and ImageNet-1K, respectively. we know that DiT used 7M iterations to train ImageNet in the original work Peebles & Xie (2023), which 1) is not the amount of resource usage that can be achieved in general research. Meanwhile, 2) the focus of this paper is on the acceleration effect of training, and a direct comparison of the ultimate convergence stage is very unnecessary for the research topic of this paper.
A.2 Instructions for code
Code is provided in the link at the very beginning of the main paper.
A.3 Architecture and Training Recipe.
We utilize Unet and DiT as our base architecture in the diffusion model. pre-trained VAE which loads checkpoints from huggingface is employed to be latent encoder. Following Unet implementation from LDM and DIT from official implementation, we provide the architecture detail in Tab. 7. We provide our basic training recipe and evaluation setting with specific details in Tab. 8.
@@ A.2 Instructions for code
Table 7: Architecture detail of Unet and DiT on MetFaces, FFHQ, and ImageNet.
@@ A.4 Datasets
A.5 Detail of MDT + SpeeD Experiment
@@
Following the official implementation of MDT, We utilize DiT-S/2 and MDT-S/2 as our base architecture, whose total block number both is 12 and the number of decoder layers in MDT is $2$. We employ the AdamW (Loshchilov & Hutter, 2017) optimizer with constant learning rate 1e-4 using $256$ batch size without weight decay on class-conditional ImageNet with an image resolution of $256^{2}$. We perform training on the class-conditional ImageNet dataset with images of resolution 256x256. The diffusion models are trained for a total of 1000K iterations, utilizing a mask ratio of 0.3.
A.6 Detail of FDM + SpeeD Experiment
FDM add the momentum to the forward diffusion process with a scale that control the weight of momentum for faster convergence to the target distribution. Following official implementation, we train diffusion models of EDM and FDM. We retrain these official network architecture which is U-Net with positional time embedding with dropout rate 0.13 in training. We adopt Adam optimizer with learning rate 1e-3 and batch size 512 to train each model by a total of 200 million images of $32^{2}$ CIFAR-10 dataset. During training, we adopt a learning rate ramp-up duration of 10 Mimgs and set the EMA half-life as 0.5 Mimgs. For evaluation, EMA models generate 50K images using EDM sampler based on Heun’s $2^{nd}$ order method (Süli & Mayers, 2003).
A.7 Text-to-Image Experiment Detail
@@
A.8 Theoretical Analysis
A.8.1 Notations
@@
A.8.2 Auxiliary Lemma and Core Theorem
@@
In DDPM (Ho et al., 2020), using a simple equivariant series $\{\beta_{t}\}_{t\in[T]}$ to simplify the complex cumulative products $\{\bar{\alpha}_{t}\}_{t\in[T]}$, we obtain the following auxiliary upper bound of $\bar{alpha}_{t}$.
@@ A.8.4 Proof of Lemma 1
A.8.5 Proof of Theorem 1
Before the proof of the theorem, we note that the samples $x_{t}|x_{0}\sim\mathcal{N}(\mu_{t},\mathbf{\sigma}_{t})$ have the following bounds with Lemma 1:
@@ Appendix B More Experiment Results
Super resolution with SpeeD. We employ SpeeD to super-resolution image generation on 512 $\times$ 512 MetFaces compared with vanilla DiT. We train DiT-XL/2 for 100K training iterations and compare the FID score at 50K, 100K training iterations. The batch size is $32$ for saving the GPU memory. As shown in 9, SpeeD obtain better performance than vanilla DiT at same training iterations on $512^{2}$ MetFaces dataset. It indicates that SpeeD can achieve training acceleration on super-resolution tasks.
B.1 Additional Experiments
B.1.1 100K iterations cut on FFHQ
A cut of training process comparison on FFHQ dataset are shown in Tab. 10.
Schedules linear quadratic cosine
DiT-XL/2 (Peebles & Xie, 2023) 7.8 7.6 8.0
ANT-UW (Go et al., 2023) 6.3 7.1 8.3
TskDif-Cur (Park et al., 2024b) 6.7 6.7 7.2
P2 (Choi et al., 2022) 10.1 10.2 12.3
MinSNR (Hang et al., 2023) 7.6 7.6 11.2
Log-Normal (Esser et al., 2024) 15.6 14.9 19.2
CLTS (Xu et al., 2024) 7.9 7.7 9.1
SpeeD(ours) 5.8 5.8 6.4
Table 10: 100K iterations cut on FFHQ.
B.1.2 Guidance Scale
Ablation experiments about guidance scale in related conditional generation task are shown in Tab. 11.
Guidance 0.75 1 1.5 2 3
t2i DiT-XL/2 20.1 22.5 27.4 31.0 36.2
t2i SpeeD 19.4 20.8 25.3 28.8 34.8
ccg EDM 1.8 1.8 1.9 1.9 2.5
ccg SpeeD on EDM 1.7 1.7 1.8 1.8 2.3
Table 11: Text-to-image (t2i) task on MS-COCO at 400K iteration and EDM on CIFAR-10 with 200M training images for class-conditional generation (ccg).
B.1.3 DiT Architecture
Ablation experiments about DiTs of different architectures are shown in Tab. 12.
S/2 B/2 XL/2
DiTs 18.1 12.9 7.8
SpeeD 15.3 10.8 5.8
Table 12: Different size of DiT with 100K iterations on FFHQ.
B.1.4 Detailed Training Process
The detailed training process on FFHQ through 100K iterations are shown in Tab. 13.
Iterations (K) 10 20 30 40 50 60 70 80 90 100
DiT-XL/2 356.1 335.3 165.2 35.8 12.9 11.9 10.5 9.6 8.7 7.8
SpeeD 322.1 320.0 91.8 19.8 9.9 7.6 7.1 6.6 6.2 5.8
Table 13: Details about training to 100K on FFHQ.
Appendix C More Related Works
@@
We discuss other works related to SpeeD, including Text to Image and Video generation. Another point to mention is that we learn from InfoBatch (Qin et al., 2023) in writing.
@@
Text to image generation with diffusion models Text-to-image generation has emerged as a hotly contested and rapidly evolving field in recent years, with an explosion of related industrial products springing up (Saharia et al., 2022; Rombach et al., 2022; Betker et al., 2023; Chen et al., 2023a; Esser et al., 2024). Convert textual descriptions into corresponding visual content, models not only learn to synthesize image content but also ensuring alignment with the accompanying textual descriptions. To better align images with textual prompt guidance, previous work has primarily focused on enhancements in several schemes including strengthening the capacity of text encoder (Raffel et al., 2020; Radford et al., 2021) improving the condition plugin module in diffusion model (Zhang et al., 2023b), improving data quality (Betker et al., 2023).
@@
Video generation with diffusion models. As diffusion models achieve tremendous success in image generation, video generation has also experienced significant breakthroughs, marking the field’s evolution and growth. Inspired by image diffusion, pioneering works such as RVD Yang et al. (2022) and VDM Ho et al. (2022b) explore video generation using diffusion methods. Utilizing temporal attention and latent modeling mechanisms, video diffusion has advanced in terms of generation quality, controllability, and efficiency Ho et al. (2022a); Singer et al. (2022); Zhou et al. (2022); He et al. (2022); Zhang et al. (2023a); Guo et al. (2023); Wu et al. (2022); Wang et al. (2023). Notably, Stable Video Diffusion Blattmann et al. (2023) and Sora Brooks et al. (2024) achieve some of the most appealing results in the field.
@@
Other diffusion acceleration works To achieve better results with fewer NFE steps, Consistency Models Song et al. (2023) and Consistency Trajectory Models Kim et al. (2023) employ consistency loss and novel training methods. Rectified Flow Liu et al. (2022), followed by Instaflow Liu et al. (2023), introduces a new perspective to obtain straight ODE paths with enhanced noise schedule and improved prediction targets, together with the reflow operation.